OpenStudy (anonymous):

Part A: Solve F = alt=9/5 C + 32 for C. (4 points) Part B: Determine the value of C when F = 10 °F. (2 points) Part C: Solve -np - 70 > 40 for n. Show your work. (4 points)

OpenStudy (anonymous):

Part A. The first thing you need to do is isolate C re work the original equation F=9/5C+32 so that C is one the other side of the = by itself. First subtract 9/5C from both sides F-9/5C=32 Then subtract F from both sides -9/5C=-F+32 Then you have to isolate C by multiplying both sides by -5/9 C=5/9F-283/9 Part B. Plug in 10 for F C=5/9*10-283/9 C=50/9-283/9 C=233/9 or C=25.89 Part C. Solve -np-70>40 for n Isolate n add 70 to both sides -np>110 then divide by -p n>-110/p

OpenStudy (anonymous):

@lilitherien thanks but I already finished my work :D

OpenStudy (denisaboichuk):

(